What is the short-run profit maximizing level of output for a firm with short-run total cost given by TC = 50 + 2q + 2q^2 that operates in a perfectly competitive industry where the short-run market demand and supply curves are given by QD = 1410 – 40P and QS = -390 + 20P?
Answer
a.
1800
b.
1130
c.
7
d.
30
1
Expert's answer
2013-07-31T09:42:55-0400
TC = 50 + 2q + 2q^2
QD = 1410 – 40P
QS = -390 + 20P
Answer:
Qd = Qs, so
1410 - 40P = -390 + 20P
60P = 1800
P = 30,
Q = 1410 - 40*30 = 210
For profit maximization MR = MC = P = 30
MC = TC' = 4Q + 2 = 30
4Q = 28
Q = 7
So, the short-run profit maximizing level of output for a firm is 7 units.
Right answer:
c. 7